# Domain Consistency Representation Learning for Lifelong Person Re-Identification

### v2

@@ Abstract
Lifelong person re-identification (LReID) exhibits a contradictory relationship between intra-domain discrimination and inter-domain gaps when learning from continuous data. Intra-domain discrimination focuses on individual nuances (i.e., clothing type, accessories, etc.), while inter-domain gaps emphasize domain consistency. Achieving a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps is a crucial challenge for improving LReID performance. Most existing methods strive to reduce inter-domain gaps through knowledge distillation to maintain domain consistency. However, they often ignore intra-domain discrimination. To address this challenge, we propose a novel domain consistency representation learning (DCR) model that explores global and attribute-wise representations as a bridge to balance intra-domain discrimination and inter-domain gaps. At the intra-domain level, we explore the complementary relationship between global and attribute-wise representations to improve discrimination among similar identities. Excessive learning intra-domain discrimination can lead to catastrophic forgetting. We further develop an attribute-oriented anti-forgetting (AF) strategy that explores attribute-wise representations to enhance inter-domain consistency, and propose a knowledge consolidation (KC) strategy to facilitate knowledge transfer. Extensive experiments show that our DCR model achieves superior performance compared to state-of-the-art LReID methods. Our code will be available soon.
@@
Lifelong learning, person re-identification, domain consistency representations, attribute and text guided representations.I Introduction
@@
Person re-identification (ReID) aims to retrieve the same individual across non-overlapping cameras in a large-scale database, and has achieved significant progress using uni-modal architectures such as convolutional neural networks (CNN) [3, 4] or vision transformers (ViT) [5, 6, 7]. However, when ReID models are applied to continuous datasets collected by dynamic monitoring systems, they exhibit notable performance limitations. Thus recent works have focused more on the practical problem of lifelong person re-identification (LReID), which involves learning from streaming data and maintaining strong performance across all data.
@@
At present, lifelong person re-identification (LReID) suffers from the challenge of balancing the anti-forgetting of old knowledge and learning new knowledge. Specifically, there are two main issues to solve this challenge. 1) Intra-domain discrimination. Each identity may exhibit subtle nuances of individual information (i.e., clothing type, accessories, haircut, etc.) and lead to severe distribution overlapping. Learning discriminative representations of individuals are effective for distinguish identity information. 2) Inter-domain gaps. The dataset of each task is collected in different illumination and background, leading to inter-domain gaps. Bridging intra-domain gaps are significant for mitigating catastrophic forgetting in LReID.
@@
To address these issues, we aim to learn consistency representations that capture individual nuances in intra-domain and inter-domain consistency in LReID, striking a balance between maximizing intra-domain discrimination and minimizing inter-domain gaps. Knowledge distillation-based approaches [2, 8, 9, 10] ensure distribution consistency between the previous and current datasets to alleviate catastrophic forgetting. However, these approaches impose strict constraints and ignore intra-domain discrimination, [11, 12, 13], as outlined in Figure 1(a). While LReID models significantly improve intra-domain discrimination for the current step, they inevitably damage inter-domain consistency, leading to catastrophic forgetting. Thus, we propose consistency representations as a bridge to achieve a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps, improving the anti-forgetting and generalization capabilities of the LReID model, as illustrated in Figure 1(b).
@@
Specifically, we propose a novel domain consistency representation learning (DCR) model, which first explores attribute and text information to enhance LReID performance. Unlike methods [14, 15, 16], we develop consistency representations including global and attribute-wise representations to capture individual nuances in intra-domain and inter-domain consistency in LReID. We design an attribute-text generator (ATG) to dynamically generate text-image pairs for each instance, which are then fed into a text-guided aggregation (TGA) network to improve the global representation capability, effectively distinguishing identities in LReID. In addition, the attributes of each instance guide an attribute compensation (ACN) network to generate attribute-wise representations focusing on specific regional information of identities. We consider that attributes can ensure reliability by setting higher thresholds across datasets. Therefore, the generated attribute-wise representations and text for each instance are considered reliable in our model.
@@
In summary, we explore global representations and attribute-wise representations to strike a balance between maximizing identity-discriminative information of intra-domain and minimizing inter-domain gaps. At the intra-domain discrimination level, global representations capture whole-body information, while attribute-wise representations focus on specific regional information. When whole-body appearances or attribute-related information are very similar across identities, we combine global and attribute-wise representations to distinguish among similar identities, maximizing intra-domain discrimination. Perfect learning intra-domain discrimination can lead to catasttophic forgetting. We further develop an attribute-oriented anti-forgetting (AF) strategy that explore attribute-wise representations for bridging inter-domain gaps across continuous datasets. Additionally, knowledge consolidation (KC) is proposed to enable knowledge transfer, improving generalization capabilities. Our contributions are as follows:
@@
I-A Lifelong Person Re-Identification
Lifelong Person Re-Identification (LReID) faces a formidable challenge, aiming to address the evolving nature of person identification across various scenarios. Some works [17, 12] are proposed to tackle the issue of adapting ReID models over time while retaining knowledge gained from previous distribution. Generally, Pu et al. [17] proposed learnable knowledge graphs that adaptively facilitate the mutual exchange of new and old knowledge, thus achieving knowledge accumulation. Some works [8, 1, 14, 9] aim to extract rich and discriminative representation, mitigating the risk of knowledge forgetting. Pu et al. [8] proposed a meta reconciliation normalization (MRN) for mining meta-knowledge shared across different domains. Meanwhile, ConRFL [9] maintains learnable and consistent features across all seen domains, which improves the discrimination and adaptation ability of the LReID model. In addition, some methods [17, 2, 1] mitigate catastrophic forgetting and enhance model accuracy by using rehearsal-based strategies with images stored from previous tasks. These approaches strive to reduce inter-domain gaps, ensuring distribution consistency across datasets to mitigate catastrophic forgetting. However, this strategy employs strict constraints and ignores intra-domain discrimination, limiting LReID model’s performance to learn new knowledge. In this paper, we propose consistency representations as a bridge to achieve a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps for improving the anti-forgetting and generalization capabilities of the LReID model.
I-B Vision-Language for Person Re-Identification
@@
The vision-language learning paradigms [18, 19, 20] have gained considerable popularity in recent years. Contrastive Language-Image Pre-training (CLIP) [21], establishes a connection between natural language and visual content through the similarity constraint of image-text pair. CLIP has been applied to multiple person re-identification tasks [22, 23], including text-to-image, text-based single-modality, text-based cross modality. Text-to-image methods [24, 25, 26] aims to retrieve the target person based on a textual query. Text-based single-modality works [5, 6] leverage text descriptions to generate robust visual features or to integrate the beneficial features of text and images for the person category. Text-based cross modality methods [27] employ text descriptions to alleviate visible-infrared modality gaps. Text infromation generated by prompt learning and text inversion, providing insufficient text descriptions of each identity. In this paper, we dynamically generate text-image pairs from single image to capture fine-grained global representations based on the CLIP model for improving inter-domain discrimination.
I-C Pedestrian Attribute Recognition
@@
Pedestrian attribute recognition aims to assign a set of attributes (Gender, Bag, Short/Long sleeve, and etc.) to a visual representation of a pedestrian. Deep learning-based researches [28] automatically learn hierarchical features from raw images, improving recognition accuracy. Multi-task learning approaches [29, 30, 31] leverage additional contextual information of across tasks, such as pedestrian detection or pose estimation, to significantly improve attribute recognition. Part-based methods [32, 33] divide the pedestrian image into several parts or regions, providing more accurate attribute localization. At present, the above methods have achieved significant success in improving the accuracy of attribute recognition. We are the first to explore the application of attributes to LReID from the following two aspects. 1) Attributes are converted into text descriptions for each image to enhance global representation capabilities. 2) Attributes are transformed into attribute-wise representations by specific networks to maximize intra-domain discrimination and minimize intra-domain gaps.
II Proposed Method
II-A Preliminary: Overview of Method
@@
The overview of our DCR model to achieve a trade-off between maximizing intra-domian discrimination and minimizing inter-domain gaps, is depicted as Figure 2. The DCR model learns the old model and new model from (t-1)-th and t-th steps, where is inherited from . and with three branches of attribute-text generator (ATG), text-guided aggregation network (TGA) and attribute compensation network (ACN). and serve as classifier heads for the old and new models, providing logits of each instance for recognition. Additionally, we define that consecutive person datasets are collected from different environments, and establish a memory buffer to store a limited number of samples from each previous ReID dataset. Given an image , we forward it to and is as follows:
II-B Attribute-Text Generator
Due to the lack of text-image pairs in ReID datasets, we propose an attribute-text generator (ATG) to dynamically generate corresponding text descriptions for each instance. Specifically, we first introduce an attribute recognition model pre-trained on the PA100K dataset [35] to generate attribute categories (i.e., female, backpack, short/long sleeve, and etc.), which are then converted into text descriptions for each instance using a specific template. This template adds modifiers (in black font) to each attribute (in a different color font) to create a complete sentence describing an instance, as shown in Figure 2. Although, attributes can vary significantly across datasets, we consider that text descriptions can be made reliable by setting a higher threshold (Confidence threshold=0.80) to ensure classification accuracy of attribute recognition network.
II-C Text-Guided Aggregation Netwrok
@@
We propose a text-guided aggregation network (TGA) to explore global representations for each identity and knowledge transfer, as shown in Figure 2 (TGA). The TGA includes a CLIP model and a parallel fusion module (PFM). Note that the text encoder is frozen in our DCR model.
@@
Parallel Fusion Module. By attribute-text generator obtain text-image pairs, we employ CLIP with text encoder and image encoder to extract text and image embedding, respectively. Unlike CLIP [21], we introduce multiple [CLS] embeddings into the image encoder input sequence to capture multiple global representations from different perspectives. To obtain fine-grained global representations for improving the performance of the LReID model, we propose a parallel fusion module (PFM) to explicitly explore the interactions between image embeddings and text embeddings, as shown in Figure 2 (PFM). Firstly, we leverage text embedding as query and image embedding [, , , , , ] as key and value to implement operation with cross-attention, drop, and layer normalization, getting text-wise representations. Similarly, in another fusion branch, image-wise representations are obtained. Finally, image-wise and text-wise representations perform concatenation and MLP operations to obtain global representations , focusing on whole body information. We force multiple global reprsentatios at the current step to learn more discriminative information by orthogonal loss to minimize the overlapping elements. The orthogonal loss can be formulated as:
@@
II-D Attribute Compensation Network
We force attributes to guide attribute compensation network (ACN) for learning attribute-wise representations. The ACN consists of an attribute decoder and an attribute matching component, as illustrated in Figure 2 (ACN).
Attribute Decoder. Enabling attributes to better adapt across datasets, we define multiple learnable attribute semantic information to learn discriminative information. The attributes undergo a linear layer to increase its dimensions, and then multiplies with the text-image global representation to output . Attribute semantic information as queries , as keys and values are input into attribute decoder, which outputs the attribute features . The attribute decoder employs six transformer blocks (TBlock) referenced from [36].
Attribute Matching. The attribute features learn multiple discriminative local infromation of individuals. However, it is unclear which attribute features correspond to specific body parts. Thus, we propose an attribute matching (AM) component to associate attribute features and global representations . The core objective is to find the most similar global representations from different perspectives and local attribute features , and then add the them with the highest similarity. Specifically, attribute-wise representations is formulated as:
@@ II-D Attribute Compensation Network
II-E Attribute-oriented Anti-Forgetting
We develop an attribute-oriented anti-forgetting (AF) strategy to explore attribute-wise representations for aligning the distributions of the old and new models, as shown in Figure 2 (AF). Specifically, the new model adapts to new information but may forget old knowledge from the previous dataset, while the old model retains old knowledge. To better preserve old knowledge, we leverage attribute-wise representations as a bridge to optimize both the old and new models using samples from the memory buffer. This strategy achieves domain consistency and minimize inter-domain gaps, alleviating the forgetting of old knowledge, and is calculated as follows:
where refers to a kullback-leibler divergence, and is a hyperparameter called temperature [37].

### v4

@@ Abstract
Lifelong person re-identification (LReID) exhibits a contradictory relationship between intra-domain discrimination and inter-domain gaps when learning from continuous data. Intra-domain discrimination focuses on individual nuances (i.e., clothing type, accessories, etc.), while inter-domain gaps emphasize domain consistency. Achieving a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps is a crucial challenge for improving LReID performance. Most existing methods strive to reduce inter-domain gaps through knowledge distillation to maintain domain consistency. However, they often ignore intra-domain discrimination. To address this challenge, we propose a novel domain consistency representation learning (DCR) model that explores global and attribute-wise representations as a bridge to balance intra-domain discrimination and inter-domain gaps. At the intra-domain level, we explore the complementary relationship between global and attribute-wise representations to improve discrimination among similar identities. Excessive learning intra-domain discrimination can lead to catastrophic forgetting. We further develop an attribute-oriented anti-forgetting (AF) strategy that explores attribute-wise representations to enhance inter-domain consistency, and propose a knowledge consolidation (KC) strategy to facilitate knowledge transfer. Extensive experiments show that our DCR achieves superior performance compared to state-of-the-art LReID methods. Our code is available at https://github.com/LiuShiBen/DCR.
@@
Lifelong person re-identification, attribute-text generator, text-image aggregation, domain consistency representation.I Introduction
@@
Person re-identification (ReID) aims to retrieve the same individual across multiple cameras in a large-scale database by using uni-modal architectures such as convolutional neural networks (CNN) [1, 2, 3] or vision transformers (ViT) [4, 5, 6]. However, when ReID models are applied to continuous datasets collected by video-based monitoring systems [7, 8], they exhibit notable performance limitations. As a result, recent works have focused on the practical problem of lifelong person identification (LReID), which maintains strong performance with continuously updated data streams.
@@
At present, lifelong person re-identification (LReID) suffers from the challenge of balancing the anti-forgetting of old knowledge and learning new knowledge. Specifically, there are two main issues to solve this challenge. 1) Intra-domain discrimination. Each identity may exhibit subtle nuances of individual information (i.e., clothing type, accessories, haircut, etc.) and lead to severe distribution overlapping. Learning discriminative representations of individuals are effective for distinguish identity information. 2) Inter-domain gaps. Each Domain is collected in different illumination and background, leading to inter-domain gaps. Bridging intra-domain gaps are significant for mitigating catastrophic forgetting in LReID.
@@
To address these issues, we aim to learn domain consistency representations that capture individual nuances in intra-domain and inter-domain consistency in LReID. Knowledge distillation-based approaches [10, 11, 12] ensure distribution consistency between the previous and current domain to alleviate catastrophic forgetting. However, these approaches impose strict constraints and ignore intra-domain discrimination, [13, 14, 15], as outlined in Fig. 1(a). While LReID models significantly improve intra-domain discrimination for the current step, they inevitably damage inter-domain consistency, leading to catastrophic forgetting. Thus, we explore global and attribute-wise representations to strike a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps, improving the anti-forgetting and generalization capabilities of the LReID model, as illustrated in Fig. 1(b).
@@
Specifically, we propose a novel domain consistency representation learning (DCR) model that first explores attribute and text information to enhance LReID performance. Unlike methods [16, 17, 18], we develop domain consistency representations including global and attribute-wise representations to capture individual nuances in intra-domain and inter-domain consistency in LReID. We design an attribute-text generator (ATG) to dynamically generate text-image pairs for each instance, which are then fed into a text-guided aggregation (TGA) network to enhance the global representation capability, effectively distinguishing identities in LReID. Furthermore, the attributes of each instance guide an attribute compensation (ACN) network to generate attribute-wise representations that focus on specific regional information about identities. We consider that attributes can enhance reliability by setting higher thresholds across domains. Therefore, the generated attribute-wise representations and text for each instance are considered reliable in our model.
@@
In summary, we aim to strike a balance between maximizing intra-domain identity-discriminative information and minimizing inter-domain gaps by exploring global and attribute-wise representations. At the intra-domain level, global representations capture whole-body information, while attribute-wise representations focus on specific regional information. When whole-body appearances or attribute-related information are similar across identities, we combine global and attribute-wise representations to distinguish among similar identities, maximizing intra-domain discrimination. While the model adapts well to new information, it often struggles with catastrophic forgetting. To preserve old knowledge, we develop an attribute-oriented anti-forgetting (AF) strategy that explores attribute-wise representations to bridge inter-domain gaps across continuous datasets. Meanwhile, striking a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps is crucial for improving the performance of LReID models. Knowledge consolidation (KC) is proposed to facilitate knowledge transfer and enhance generalization capabilities, which consists of alignment and logit-level distillation mechanisms. The alignment mechanism explores global representations of knowledge transfer from the current dataset. The logit-level distillation mechanism enhances the extraction of identity information, thereby improving the model’s ability to consolidate knowledge. Our contributions are as follows:
@@
II Related work
II-A Lifelong Person Re-Identification
Lifelong Person Re-Identification (LReID) aims to balance intra-domain discrimination with minimizing inter-domain gaps in continuously updated datasets across scenarios, improving the model’s anti-forgetting and generalization capabilities. LReID methods can be divided into two categories. 1) Knowledge distillation-based methods [12, 19, 20, 21] utilize metric strategies to achieve domain-consistent alignment between the old model with learned knowledge distribution and the new model that adaptively learns new knowledge. 2) Exemplar-based methods [22, 10, 9] achieve a distribution balance between old and new samples to prevent catastrophic forgetting by forming a memory buffer to select the limited samples from some identities. These methods strive to reduce inter-domain gaps and ensure consistency across domains to prevent catastrophic forgetting. However, they ignore intra-domain identity discrimination and lack consistency optimization within the inter-domain, limiting the LReID model’s performance in learning new knowledge. In this paper, we explore domain consistency representations as a bridge to achieve a trade-off between maximizing intra-domain discrimination and minimizing inter-domain gaps for enhancing the anti-forgetting and generalization capabilities of the LReID model.
II-B Vision-Language for Person Re-Identification
@@
Vision-language learning paradigms [23, 24] have gained widespread popularity in recent years. Contrastive Language-Image Pre-training (CLIP) [25], establishes a connection between natural language and visual content through the similarity constraint of image-text pairs. CLIP has been applied to multiple person re-identification tasks [26, 27, 28], including text-to-image, text-based single-modality, and text-based cross-modality. Text-to-image methods [29, 28, 30] aim to retrieve the target person based on a textual query. Text-based single-modality works [31, 27, 5] leverage text descriptions to generate robust visual features or integrate the beneficial features of text and images for the person category. Text-based cross-modality methods [32] utilize text descriptions to reduce visible-infrared modality gaps. Providing insufficient text descriptions of each identity, due to prompt learning and text inversion. In this paper, we dynamically generate text-image pairs from single images to capture fine-grained global representations based on the CLIP model for improving model performance capability in terms of inter-domain and intra-domain.
II-C Pedestrian Attribute Recognition
@@
Pedestrian attribute recognition aims to assign a set of attributes (Gender, Bag, Short/Long sleeve, and etc.) to a visual representation of a pedestrian based on their attributes. Deep learning-based research [33] automatically learns hierarchical features from raw images, improving recognition accuracy. Multi-task learning methods [34, 35, 36] leverage additional contextual information from multiple domains, such as pedestrian detection or pose estimation, to significantly improve attribute recognition. Part-based methods [37, 38] divide the pedestrian image into several parts or regions, providing more accurate localization. Currently, these methods have achieved significant success in improving the accuracy of attribute recognition. We are the first to explore the application of attributes to LReID from two perspectives. 1) Attributes are converted into text descriptions for each image to enhance global representation capabilities. 2) The attributes are transformed into attribute-wise representations by specific networks to maximize intra-domain discrimination and minimize intra-domain gaps.
III Proposed Method
III-A Preliminary: Overview of Method
@@
The overview of our DCR to achieve a trade-off between maximizing intra-domain discrimination and inter-domain gaps is shown in Fig. 2. The DCR model learns the old model and new model from (t-1)-th and t-th steps, where is inherited from . and with three parts of attribute-text generator (ATG), text-guided aggregation network (TGA), and attribute compensation network (ACN). and serve as classifier heads for the old and new models, providing logits of each instance for recognition. Additionally, we define that consecutive person datasets are collected from different environments, and establish a memory buffer to store a limited number of samples from each previous ReID dataset. Given an image , we forward it to and is as follows:
Where and are global and attribute-wise representations, respectively.
III-B Attribute-Text Generator
Due to the lack of text-image pairs in ReID datasets, we propose an attribute-text generator (ATG) to dynamically generate text-image pairs for each instance, as illustrated in Fig. 2. Firstly, we employ a pre-trained attribute recognition model [33] (trained on the PA100K dataset [39]) to generate attribute categories. These categories are organized into four major types (e.g., overall, upper body, lower body, and decoration) with 12 subcategories, as shown in TABLE I. Then, attribute categories are converted into text descriptions for each instance using a predefined template. Finally, the generated text-image pairs are fed into the text-guided aggregation network (TGA) for further processing. In this predefined template (Fig. 2), black font denotes fixed modifiers, while red, blue, orange, and green fonts represent the four major types, respectively. The template dynamically replaces the colored placeholders with the corresponding predicted attributes. To ensure robustness across domains, we set a high confidence threshold (0.80) for attribute classification, thereby maintaining the reliability of text descriptions despite potential domain variations.
| Type | Category | 0 | 1 |
| Overall | Sex | man | woman |
| Upper body | Short sleeved top | No | Yes |
| Long sleeved top | No | Yes | |
| Long coat | No | Yes | |
| Lower body | Trousers | No | Yes |
| Shorts | No | Yes | |
| Skirt | No | Yes | |
| Decoration | Hat | No | Yes |
| Glasses | No | Yes | |
| Handbag | No | Yes | |
| Sholder bag | No | Yes | |
| Backpack | No | Yes |
III-C Text-Guided Aggregation Network
@@
To enhance robust representation capabilities for each instance, we propose a text-guided aggregation network (TGA) to explore the global representation for distinguishing identity information in LReID, as shown in Fig. 2 (TGA). The TGA consists of a Contrastive Language–Image Pretraining(CLIP) model and a parallel fusion module (PFM). Note that the text encoder is frozen in our DCR model.
III-C1 CLIP Model
The CLIP model is a vision-language model that learns multi-modal representations by aligning images and text in a shared embedding space. The model employs a contrastive learning objective to maximize the similarity between an image and its corresponding text description. In our work, we adapt the CLIP model, which includes an image encoder and a text encoder, to generate both image and text embeddings. However, unlike the original CLIP model [25], our model introduces multiple [CLS] embeddings into the image encoder’s input sequence. This modification enables the model to capture diverse representations from different views for each instance, thereby enhancing its ability to distinguish fine-grained identity information. Specifically, the clip model generate text embedding and image embedding [, , , , , ] for extracting rich text and vision information. image embedding consists of multiple [CLS] embeddings [, , ] for multi-view representations and patch embedding [, , ] for local feature extraction.
III-C2 Parallel Fusion Module
@@
To improve the performance of the LReID model, we propose a parallel fusion module (PFM) to explicitly explore the interactions between image and text embeddings, as shown in Fig. 2 (PFM). Firstly, we leverage text embedding as query and image embedding [, , , , , ] as key and value to implement operation with cross-attention, drop, and layer normalization, getting text-wise representations. Similarly, in another fusion branch, image-wise representations are obtained. Finally, image-wise and text-wise representations perform concatenation and MLP operations to obtain global representations , focusing on whole body information. We force multiple global representations at the current step to learn more discriminative information by orthogonal loss to minimize the overlapping elements. The orthogonal loss can be expressed as:
@@
III-D Attribute Compensation Network
We force attributes to guide the attribute compensation network (ACN) for learning attribute-wise representations. The ACN consists of an attribute decoder and an attribute matching component, as illustrated in Fig. 2 (ACN).
III-D1 Attribute Decoder
Enabling attributes to better adapt across domains, we define multiple learnable attributes semantic information to learn discriminative information. The attributes undergo a linear layer to increase their dimensions and are then multiplied with the text-image global representation to output . Attribute semantic information as queries , as keys and values are input into the attribute decoder, which outputs the attribute features . The attribute decoder utilizes six transformer blocks (TBlock) referenced from [40].
III-D2 Attribute Matching
The attribute features focus on multiple discriminative local details of individuals. However, it remains unclear which views make these attribute features more prominent in the global representations. We propose an Attribute Matching (AM) mechanism to select the optimal alignment between attribute features and global representations, thereby identifying the best view for the global representation and improving the representational ability of the attribute features. Specifically, we first calculate the similarity between the attribute features and global representations from different views. Based on this matching process, we determine the index of the most similar global representation as follows.
where represent cosine similarity. The operation selects both the maximum similarity value and its corresponding index. We identify the most similar index between each attribute feature and the global representations , and then combine the attribute feature and with its corresponding global representation (where is the matched index) to obtain the attribute-wise representations . This process is defined as:
@@ III-D2 Attribute Matching
III-E Attribute-oriented Anti-Forgetting
We develop an attribute-oriented anti-forgetting (AF) strategy to explore attribute-wise representations that align the distributions of the old and new models, as shown in Fig. 2 (AF). The new model can adapt to new information but may forget old knowledge from the previous datasets, while the old model retains old knowledge. To preserve old knowledge, we leverage attribute-wise representations as a bridge to optimize both the old and new models by using samples from the memory buffer. This strategy achieves domain consistency and minimizes inter-domain gaps, alleviating the forgetting of old knowledge, and is calculated as follows:
Where is a kullback-leibler divergence, and represents a hyper-parameter called temperature [41].
III-F Knowledge Consolidation
Maximizing intra-domain discrimination and minimizing inter-domain gaps are in a contradictory relationship. Therefore, achieving a balance between them is crucial for improving the performance of LReID models. Thus, we propose a knowledge consolidation (KC) strategy that leverages global representations for knowledge transfer between old and new models. This includes alignment mechanism and logit-level distillation mechanism.
Maintaining distribution consistency between the old and new models for previous datasets can limit the model’s ability to learn new knowledge. Therefore, we propose an alignment mechanism to explore global representations of knowledge transfer from the current dataset, as follows:
We further introduce a logit-level distillation mechanism to enhance the extraction of identity information shared between the old and new models, further improving the model’s knowledge consolidation ability. This is represented as follows:
The knowledge consolidation loss is defined as:
The total loss function is formulated as:
| Type | Datasets | Scale | Train IDs | Test IDs | |
| Seen | Market1501 [42] | large | 500(751) | 750 | |
| CUHK-SYSU [43] | mid | 500(942) | 2900 | ||
| DukeMTMC [44] | large | 500(702) | 1110 | ||
| MSMT17V2 [45] | large | 500(1041) | 3060 | ||
| CUHK03 [46] | mid | 500(700) | 700 | ||
| Unseen | VIPeR [47] | small |
316 | |
| GRID [48] | small |
126 | ||
| CUHK02 [49] | mid |
239 | ||
| OccDuke [50] | large |
1100 | ||
| OccREID [51] | mid |
200 | ||
| PRID2011 [52] | small |
649 |